Circular No. 6870

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

S/1997 U 1 Precovery and recovery observations of this satellite are reported:

UT  α_{2000}  δ_{2000}  Observer
1984 June 1.5556  16^{h}40^{m}58.^{s}4  −22^{o}03′56′′  Cruikshank
2.5410  16 40 48.5  −22 03 36  _{”}
1998 Mar. 27.79254 20 57 08.27 −17 56 16.2 Smith
31.40380 20 57 41.27 −17 53 56.1 Hall

D. P. Cruikshank (Mauna Kea). Measurer B. Gladman. Images found on

plates taken with the Canada-France-Hawaii Telescope; uncertainty ±3′′. R. M. Smith, J. B. Jones and D. Windridge (Anglo-Australian Observa-

tory). 4-m reflector + CCD. Measurer B. Gladman. P. Hall and D. Graham (Cerro Tololo Interamerican Observatory). 4-m

reflector + CCD. Measurer J. J. Kavelaars.

Orbit computations by G. V. Williams, K. Aksnes and the undersigned show that the 1998 recovery observations, although some 3′ from the prediction on IAUC 6833, confirm that this satellite has a low-e orbit (see also IAUC 6780 and 6765). They also clearly demonstrate the validity of the 1984 candidate, already noted by Gladman prior to the 1998 recovery, but not previously announced.

Epoch = 1998 July 6.0 TT

T = 1998 Apr. 19.6463 TT  ω = 339.^{o}4621
e = 0.082347  Ω = 174.9928 }2000.0
q = 0.043975 AU  i = 139.6813

a = 0.047921 AU n^{o} = 0.621291 P = 579 days

1998TT α_{2000} δ_{2000} Δ r_{U} ϵ Δα′′ Δδ′′
Apr. 7 20^{h}58.^{m}61 −17^{o}50.^{′}0 20.243 0.044 65.^{o}0 329− 248−
17 20 59.76 −17 45.0 20.082 0.044 74.5 328− 215−
27 21 00.61 −17 41.2 19.913 0.044 84.1 322− 178−
May 7 21 01.14 −17 38.7 19.742 0.044 93.7 311− 138−
17 21 01.35 −17 37.6 19.573 0.044 103.3 296− 95−
27 21 01.23 −17 37.8 19.411 0.044 112.9 277− 49−
June 6 21 00.80 −17 39.4 19.261 0.045 122.6 253−  2−
16 21 00.07 −17 42.1 19.127 0.045 132.3 225− 46+
26 20 59.07 −17 45.9 19.013 0.045 142.1 194− 93+
July 6 20 57.86 −17 50.6 18.922 0.045 151.9 160− 139+
16 20 56.47 −17 56.0 18.858 0.046 161.8 124− 184+
26 20 54.96 −18 01.8 18.822 0.046 171.7  85− 226+
1998 April 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden